2.Based on the following information answer questions

Holyproducts corporations have the folowing capital structure, which it considers optimal:

Retained earnings 300,000

Bonds, 7% (at par) Br 300,000

 Preferred stock, Br.5 240,000

Common stock 360,000

 1,200,000

Additiona Informaton:

Dividends on common stock are currently Br 3 per share and are expected to grow at constant 

A.Based on the above information, what woud be the cost of the bond?

rate of 6%.

Market price of common stock is Br 40 and the preferred stock is selling at Br50.

Flotation cost on new issues of common stock is 10%.

The interest on bonds is paid annualy and the company’s tax rate is 40%.

B.What would be the cost of common stock for Hoy Products Corporation?


1
Expert's answer
2022-01-06T03:32:23-0500

(A):

Cost of bond = 7% (annual interest on bonds)


(B):

Cost of common stock for Hoy Products Corporation is:

"R_E=\frac{D_1}{P_0}+g"


"R_E=\frac{3\times0.06}{40}+0.06"


"R_E=\frac{0.18}{40}+0.06=0.0645"


Cost of common stock is 6.45%.